# Towards Immersive Mixed Reality Street Play

### v1

@@ 5.3.1. Fantasy world overlaid on the real world (23/63)
Participants onsite think this MR street play provided them with a unique opportunity to immerse themselves in a magical fantasy world that seemingly came to life in reality. Players expressed a strong sense of being transported into the realm of their dreams and fantasies, where they could embody the roles of powerful wizards and engage in long-awaited duels. One player exclaimed, ”I have been waiting for this duel all my life!” (P13) Another participant expressed her excitement:
”You have fulfilled my wildest fantasies. This is why I’m here.” (P14)
@@
The game evokes a strong emotional connection to fantasy lovers and their childhood dreams. A participant commented: ”I feel like a wizard. I think my letter from Hogwarts is finally gonna come in the mail after all these years.” (P25) Referencing the Harry Potter series, their words showcased how this game tapped into the collective imagination and allowed players to embody the roles of their favorite fictional characters. These reviews show that the seamless integration of immersive MR technology and fun gameplay contributed to a deep sense of immersion, making the fantasy feel tangible and real for the players.
5.3.2. Wand as a meditator for intuitive interaction and narrative immersion (33/65, behavior)
@@
Onsite participants praised the intuitive and smooth interaction facilitated by the combination of the wand and Apple Watch in this game. Unlike traditional handheld controllers, this design allowed for natural and immersive casting of spells in the real world. Participants quickly grasped the gameplay mechanics. In the on-site study, 24/65 participants were observed to immediately understand how to cast spells after the game started, either by listening to oral instructions or observing other players. Another 31/65 participants learned the mechanics within 30 seconds with guidance from organizers and fellow or previous players. As one remarked, ”It’s so easy to use (the wand) too, very comfortable and functional” (P15). Another participant noted, ”The strange rhythm of beating is actually quite strong” (P6).
@@
The wand also served as a crucial element in the narrative design, providing a compelling explanation for the asymmetrical perspectives between players and passersby. By assuming the role of wizards and perceiving passersby as Muggles, players could engage in the MR experience smoothly. The physical presence of the wand helped bystanders understand the ongoing MR experience, even without an MR view. ”It’s just making so much sense when you see people waving a wand. Much better than randomly stubbing your hands.” (P4)
@@
The significance of the wand was further highlighted during a demo where it was taken away from a player’s hand in the later stage of the game. The player immediately struggled to perform as she had been doing just 5 seconds earlier, expressing her confusion and requesting the wand back:
”Now I forgot how to move… Can you give me back the wand?” (P8).
This incident demonstrated how the presence of a prop can support players’ understanding and engagement, and maintain the narrative immersion.
5.3.3. Exercise and work-out (14/63)
@@
The game led to a notable increase in physical activity and exercise among players. Many participants (26/65, behavior) were observed breathing heavily after playing, indicating significant physical exertion. The level of physical activity varied depending on the game mode, with the single-player game resulting with minimal movement, the duel game encouraging more movement and dodging with the opponent, and the dragon game resulting in the most physical activity, with players running to chase the dragon or avoid fireball attacks.
@@
Participants appreciated the physical demands of the game, with some expressing surprise at the level of exertion required (”This is such a workout!”) (P16,17). Another compared it to playing a ”fierce badminton game” (P9), suggesting that MOFA could serve as an engaging alternative to traditional forms of exercise. The physical nature of the gameplay was also seen as a distinguishing factor compared to traditional PC games, as noted: ”The main difference from PC games is its physical exertion.” (P7)
5.3.4. Frustration caused by technological inadequacies (12/63)
Despite the overall positive gameplay experience, some participants expressed dissatisfaction with certain technological aspects of the MR street play MOFA. One of the main concerns was the form factor and weight of the devices used. One commented, ”This is not the form factor that will change the paradigm (for MR)” (L2). Another person who is interested in the game expresses concern about the device ”But iPhones are really heavy” (L29).
Participants also complained about the lack of body movement input and prompt feedback. ”It didn’t look like I sent it out with my own hands. the angle of my hand, the angle of my body, and the angle of my head will not affect the result” (P7). Similarly, another participant stated, ”I have no way of judging how precise I am going to be, it’s hard to go” (P8).
@@ 5.3.4. Frustration caused by technological inadequacies (12/63)
Another issue raised was the precision of the anchoring system. ”Sometimes the shield is here and there” (P3). Some other participants found the connecting technology and user flow to be cumbersome. ”I think the connecting could be a bit easier” (P18).
@@
In terms of MR imaging, the transparency of the optical see-through AR became a point of criticism, with one participant stating, ”It (AR part) is a little too transparent.” (P26, P27). The limited field of view was also a concern, ”It can only show a part of the dragon, and I feel that it’s trapped in the screen” (P11).
@@ 5.4. Environment and Location
The environment provides context for gameplay, influencing the gameplay excitement, actual play movement, and social acceptance extent.
@@ 5.4.1. Landscape acts as a vital backdrop
The surrounding environment setting can be seen as the backdrop of the game map and plays a crucial role in players’ emotional envokes. A proper backdrop of landscape allows players to fully engage with the game’s mission and narrative in a meaningful way. Environments that fit the game content can enhance the game experience. In S5, researchers organized to demonstrate and play MOFA the Dragon on a snow mountain. The environment and natural scenery gave the participants a feeling of awe and epic. The small group of players and spectators started to call the players ”heroes” and ”heroine” in the process. After the test, one of the heroine”* posted on her social media an image of her playing the game together with the text:”Today I feel like I am in a dream.” (P30)
On the other hand, the same game experience can be diminished in environments that do not align with the thematic elements of the game, for instance, attempting to play the dragon hunting game in a small, low-ceiling interior setting may feel less immersive and impactful.
5.4.2. Real world acts as the digital game terrain
@@
The design of MR environments is not solely confined to the digital realm; the physical environment serves as a ”game terrain” that influences players’ behaviors in games. In ”The Duel”, participants frequently navigate around pillars or tables for cover and shielding (P3, P4, P11, P12). The MOFA games currently do not integrate occlusion function in gameplay, meaning that obstacles in the physical world do not impede attacks; moreover, the shield and spells visually also are not affected by the occlusion in players’ views. Despite of that, players still inherently engage in these strategies, as their vision tracks opponents’ movements rather than merely relying on MR content. Participants also expressed expectations of future gameplays reacting to the physical environment, like spells bouncing off walls and grounds, or ”imagine a Charizard (Pokemon GO dragon) standing on top of my local mall” (L22), considering ”The connectivity to the building and the space will be the strong point of the game” (P12).
5.4.3. Context Decides Social Acceptance
@@
The spatial context, including the environment and location, plays a crucial role in determining social acceptance and adherence to social norms when it comes to immersive MR street play. People perceive that open spaces, such as parks, are more suitable for engaging, while crowded and busy spaces or potentially unsafe situations, like sidewalks, are considered inappropriate. One online reviewer expressed a strong objection to the video depicting people playing the MOFA games: ”The park, yes, that works but a busy street, are they just out of touch with reality?” (L13)
@@
Reviewers who favor the game envisioned the success of the game should be tied with open areas in urban parks rather than crowded areas: ”This is doing all the right things and I think I’ll get one (the game), though I think their videos are doing a disservice showing people playing in crowded areas. Show people playing in parks surrounded by people bored on picnic blankets…in Dolores Park when the weather turns nice in April.” (L16)
Certain venues also provide social acceptance to the immersive MR street play, as one commented ”If Pokemon GO had this during its hay-day the dork factor would have been socially acceptable” (L15).
@@ 6.1. Immersive MR Street Play in Public: From Dream to Reality
[…]
Visions of MR headsets transforming public street play have been widely depicted across various media—from films and promotional tech demonstrations to game trailers. While this remains a ”dream” for AR enthusiasts, most people have not yet experienced this kind of play. Through realizing this dream on deploying our open-source exploratory game probe series MOFA into the real world, we glimpse the future of IMRSP, providing an initial yet foundational understanding of the social implications of this novel form of play.
[…]
Although prior Human-Computer Interaction (HCI) research has extensively explored interactive public systems—such as large public displays (Cao et al., 2008), urban AR games like Pokémon GO (Paavilainen et al., 2017), and human-robot interactions in city settings (Pelikan et al., 2024)—the specific context of collocated, bodily mixed reality play using see-through MR headsets in actual urban streets remains underexplored. Our study directly addresses this gap by providing one of the first empirical investigations of IMRSP within authentic street environments, beyond controlled or private settings.
[…]
@@
[…]
Furthermore, our observations demonstrate how IMRSP actively reshapes everyday urban social dynamics, not merely overlaying digital content onto physical spaces but transforming streets into interactive stages for playful encounters. Echoing earlier research that identifies the street as a vibrant, inhabited site of socially organized human action (Pelikan et al., 2024; Stevens, 2007), our study found MR street play creates shared, performative experiences among players, bystanders, and passersby. The urban street environment—traditionally governed by subtle social norms and unspoken cues—becomes a dynamic arena of spectacle and interaction, actively drawing passersby into novel social engagements.
@@
Our findings indicate that IMRSP significantly reshapes social dynamics in public spaces, aligning closely with established concepts from prior research on public interactive installations. One notable parallel is the ”honeypot effect,” (Wouters et al., 2016) previously observed in interactive displays and street-level activities (Cao et al., 2008; Reeves et al., 2005a), where visible, playful activities naturally draw passersby into becoming spectators and even spontaneous participants. Similarly, during our field trials, once a small group initiated MR street play, curious bystanders consistently gathered around them. Players immersed in digital battles or performing dramatic gestures transformed ordinary streets into lively theaters of interaction. This mirrors previous findings showing how street play and public interactive technologies dynamically convert urban spaces into stages for public performance and social interaction (Stevens, 2007; Benford et al., 2013a).
@@
Beyond merely drawing attention, IMRSP created significant social contagion effects. Our observations demonstrated that players felt notably less awkward engaging in playful activities when accompanied by peers. The presence of a supportive social group reduced inhibitions, echoing findings from previous interactive installations where collective engagement validated individual participation in otherwise unconventional behaviors (Reeves et al., 2005a; Wouters et al., 2016). Additionally, bystanders frequently mirrored the players’ enthusiasm through gestures, applause, or informal engagement, reinforcing the playful atmosphere. This contagion effect underscores IMRSP’s potential to leverage social facilitation, effectively lowering barriers for hesitant individuals to engage in public MR interactions.
@@
Moreover, participants frequently embraced the public attention as an opportunity for deliberate showmanship (Reeves et al., 2005a), exaggerating movements and reactions to entertain and involve bystanders. Rather than feeling inhibited by an audience, these players became energized, further emphasizing the performative and expressive dimensions of IMRSP. This aligns with prior research on embodied interaction and performance-led HCI approaches, where technology encourages expressive and performative engagement (Dourish, 2001; Benford et al., 2013a). Such intentional performances amplified community interactions and fostered a positive feedback loop, attracting additional spectators and participants.
@@
Collectively, these findings reveal how IMRSP actively transforms public spaces into shared arenas of playful social interaction and performance. The street, traditionally governed by subtle social norms such as Goffman’s notion of ”civil inattention” (Goffman, 2008), becomes a vibrant social stage where traditional boundaries between player and audience blur. As MR street play grows more prevalent, understanding and managing these emergent social dynamics will be essential for facilitating inclusive and positive public interactions in MR-rich urban futures.
@@ 6.2. The Frictional Path to Widespread MR Street Play Adoption
6.2.1. Contrast of Social Acceptance: Onsite vs. Online Perspectives
[…]
Our study revealed a notable gap in social acceptance between those who experienced IMRSP firsthand and those who observed it remotely through recorded gameplay videos. Participants who directly engaged in MR street play consistently reported highly positive experiences, describing IMRSP as enjoyable, immersive, and socially rewarding. They appreciated the embodied interaction facilitated by MR headsets, highlighting that playing alongside friends and maintaining awareness of their physical surroundings significantly reduced feelings of isolation or awkwardness. This onsite enthusiasm aligns closely with prior research emphasizing the role of collaborative experiences in alleviating social discomfort during public interactions (Reeves et al., 2005a; Huggard et al., 2013; Benford et al., 2013a).
In sharp contrast, remote online observers, who lacked direct interaction and immersion, expressed far more ambivalence and skepticism towards IMRSP. Online commenters frequently voiced concerns about social awkwardness or embarrassment associated with observed behaviors, often characterizing public MR play as ”strange” or ”inappropriate.” Safety concerns were also prominent among these remote viewers, who worried that players might become distracted, increasing the risk of accidents or collisions in public settings. These critical perceptions resonate with earlier societal reactions to novel wearable technologies such as Google Glass, where non-users perceived the technology as intrusive or socially disruptive due to unfamiliarity and lack of direct exposure (Pelikan et al., 2024; McGrane, 1994).
This perception gap underscores a critical challenge in achieving broader public acceptance of IMRSP. Onsite players, immersed in the embodied and collaborative MR experience, quickly grasp its social and recreational value, seeing the technology as socially acceptable and enriching. Conversely, external observers, lacking firsthand immersion, primarily witness unconventional behaviors without the contextual understanding or immediate social engagement afforded by direct participation. As such, these observers often interpret IMRSP as violating established social norms or expectations, particularly regarding appropriate behavior in shared public spaces—a reaction aligned with Goffman’s insights on social norms and the management of impressions in public contexts (Goffman, 2008).
@@
6.2.2. Ethical, Safety and Privacy Concerns in Public MR Play
[…]

### v2

@@ 5.3.1. Fantasy world overlaid on the real world (23/63)
Participants on-site reported that IMRSP provided them with a unique opportunity to immerse themselves in a magical fantasy world that seemingly came to life in reality. Players expressed a strong sense of being transported into the realm of their dreams and fantasies, where they could embody the roles of powerful wizards and engage in long-awaited duels. One player (P13) exclaimed, “I have been waiting for this duel all my life!” Another participant expressed her excitement: “You have fulfilled my wildest fantasies. This is why I’m here” (P14).
@@
The game evoked a strong emotional connection to fantasy lovers and their childhood dreams. A participant (P25) commented: “I feel like a wizard. I think my letter from Hogwarts is finally gonna come in the mail after all these years.” Referencing the Harry Potter series, their words showcased how this game tapped into the collective imagination and allowed players to embody the roles of their favorite fictional characters. These reviews show that the seamless integration of immersive MR technology and fun gameplay contributed to a deep sense of immersion, making the fantasy feel tangible and real for the players.
5.3.2. Wand as a mediator for intuitive interaction and narrative immersion (33/65, behavior)
@@
On-site participants praised the intuitive and smooth interaction facilitated by the combination of the wand and Apple Watch in this game. Unlike traditional handheld controllers, this design allowed for natural and immersive casting of spells in the real world. Participants quickly grasped the gameplay mechanics. In the on-site study, 24/65 participants were observed to immediately understand how to cast spells after the game started, either by listening to oral instructions or observing other players. Another 31/65 participants learned the mechanics in less than one minute with guidance from organizers and fellow or previous players. As one participant (P15) remarked, “It’s so easy to use (the wand) too, very intuitive and functional.
@@
The wand also served as a crucial element in the narrative design, providing a compelling explanation for the asymmetrical perspectives between players and passersby. By assuming the role of wizards and perceiving passersby as muggles, players could engage in the MR experience smoothly. The physical presence of the wand helped bystanders understand the ongoing MR experience, even without an MR view. “It makes much more sense to see people waving a wand. This is far better than randomly punching the air with your hand. ” (P4).
@@
The significance of the wand was further highlighted during a demo where it was taken away from a player’s hand in the later stages of the game. The player immediately struggled to perform as she had been doing just 5 seconds earlier, expressing her confusion and requesting the wand back: “Now I forgot how to move… Can you give me back the wand?” (P8). This incident demonstrated how the presence of a prop can support players’ understanding and engagement and maintain the narrative immersion.
5.3.3. Exercise and workout (14/65)
@@
The game led to a notable increase in physical activity and exercise among players. Many participants (26/65, behavior) were observed breathing heavily after playing, indicating significant physical exertion. The level of physical activity varied depending on the game mode, with the single-player game resulting in minimal movement, the duel game encouraging more movement and dodging with the opponent, and the dragon game resulting in the most physical activity, with players running to chase the dragon or avoid fireball attacks.
@@
Participants appreciated the physical demands of the game, with some expressing surprise at the level of exertion required: “This is such a workout!” (P16, P17). Another compared it to playing a “fierce badminton game” (P9), suggesting that MOFA could serve as an engaging alternative to traditional forms of exercise. The physical nature of the gameplay was also seen as a distinguishing factor compared to traditional PC games, as P7 noted: “The main difference from PC games is its physical exertion.
5.3.4. Frustration caused by technological inadequacies (12/65)
Despite the overall positive gameplay experience, some participants expressed dissatisfaction with certain technological aspects of MOFA.
@@ 5.3.4. Frustration caused by technological inadequacies (12/65)
Participants sometimes raised issues about the precision of the anchoring system. “Sometimes the shield is here and there” (P3). Some other participants found the user flow of the co-location connection to be cumbersome. For example, “I think the connecting could be a bit easier” (P18).
@@
Participants also noted limitations of the current MR display technology, such as the transparency of the optical see-through AR and the limited field of view. For example, “it can only show a part of the dragon, and I feel that it’s trapped in the screen” (P11).
Participants also complained about the lack of prompt feedback to their body movement. “It didn’t look like I sent it out with my own hands. The angle of my hand, the angle of my body, and the angle of my head will not affect the result” (P7). Similarly, another participant (P8) stated, “I have no way of judging how precise I am going to be, it’s hard to aim.
@@ 5.4. Environment and Location
The environment provides context for gameplay, influencing excitement, movement, and the degree of social acceptance.
@@ 5.4.1. Landscape acts as a vital backdrop
The surrounding environment setting can be seen as the backdrop of the game map and plays a crucial role in players’ emotional engagement. A proper backdrop of landscape allows players to fully engage with the game’s mission and narrative in a meaningful way. Environments that fit the game content can enhance the game experience. For example, researchers once organized to demonstrate and play MOFA “the Dragon” in an alpine mountain. The natural environment and stunning scenery evoked a sense of awe in participants through the epic scale. The small group of players and spectators started to call the players “heroes” and “heroines” in the process. After the test, one of them commented that “Today I feel like I am in a dream” (P30). On the other hand, the same game experience can be diminished in environments that do not align with the thematic elements of the game. For instance, attempting to play the dragon-hunting game in a small, low-ceiling interior setting may feel less immersive and impactful.
5.4.2. The real world acts as the digital game terrain (8/65)
@@
The design of MR environments is not solely confined to the digital realm; the physical environment serves as a “game terrain” that influences players’ behaviors in games. In “The Duel”, participants frequently navigate around pillars or tables for cover and shielding (P3, P4, P11, P12). The MOFA games did not integrate real-time environment detection in gameplay, meaning that obstacles in the physical world do not impede attacks; moreover, spells and shields are not geometrically occluded in the rendering, yet players still use real objects for cover. Nevertheless, players still inherently engage in these strategies, as their vision tracks opponents’ movements rather than merely relying on MR content. Participants also expressed expectations of future gameplays reacting to the physical environment, like spells bouncing off walls and grounds, or “imagine a Charizard (Pokémon GO dragon) standing on top of my local mall” (L22), considering “The connectivity to the building and the space will be the strong point of the game” (P12).
5.4.3. Context shapes social acceptance
@@
The spatial context, including the environment and location, plays a crucial role in determining social acceptance and adherence to social norms when it comes to immersive MR street play. People perceived that open spaces, such as parks, are more suitable, while crowded and busy spaces or potentially unsafe situations, like sidewalks, are considered inappropriate. One online reviewer (L13) expressed a strong objection to the video depicting people playing the MOFA games: “The park, yes, that works, but a busy street, are they just out of touch with reality?
@@
People who favored the game envisioned that the success of the game should be tied to open areas in urban parks rather than crowded areas. For example, L16 commented that “This is doing all the right things and I think I’ll get one (the game), though I think their videos are doing a disservice showing people playing in crowded areas. Show people playing in parks surrounded by people bored on picnic blankets…in Dolores Park when the weather turns nice in April.” Certain venues also provide social acceptance to the immersive MR street play, as L15 commented “If Pokémon GO had this during its heyday, the dork factor would have been socially acceptable.
@@ 6.1. Immersive MR Street Play in Public: From Dream to Reality
Visions of MR headsets transforming public street play have been widely disseminated across various media, from films and promotional tech demonstrations to game trailers. While this remains a “dream” for tech enthusiasts, most people have not yet experienced this kind of play. Through realizing this dream of deploying our open-source exploratory game probe series MOFA into the real world, we glimpse the future of IMRSP, providing an initial yet foundational understanding of the social implications of this novel form of play.
6.1.1. Experiencing IMRSP In the Wild
Although prior HCI research has extensively explored interactive public systems—such as large public displays (Cao et al., 2008), urban AR games such as Pokémon GO (Paavilainen et al., 2017), human-robot interactions in city settings (Pelikan et al., 2024), and VR remote collaboration scenarios within constrained passenger spaces (Medeiros et al., 2024b; Ng et al., 2021; Medeiros et al., 2023)—the specific context of co-located, bodily mixed reality play using see-through MR headsets in actual urban streets remains underexplored. Our study directly addresses this gap by providing one of the first empirical investigations of IMRSP within authentic street environments, beyond controlled or private settings.
A dream you dream alone is only a dream. A dream you dream together is reality.
@@ 6.1.1. Experiencing IMRSP In the Wild
Furthermore, our observations demonstrate how IMRSP actively reshapes everyday urban social dynamics, not merely overlaying digital content onto physical spaces but transforming streets into interactive stages for playful encounters. Echoing earlier research that identifies the street as a vibrant, inhabited site of socially organized human action (Pelikan et al., 2024; Stevens, 2007), our study found that MR street play creates shared, performative experiences among players, bystanders, and passersby. The urban street environment—traditionally governed by subtle social norms and unspoken cues—becomes a dynamic arena of spectacle and interaction, actively drawing passersby into novel social engagements.
@@ 6.1.2. Augmenting Social Dynamics through IMRSP
Our findings indicate that IMRSP significantly reshapes social dynamics in public spaces, aligning closely with established concepts from prior research on public interactive installations. One notable parallel is the “honeypot effect” (Wouters et al., 2016), previously observed in interactive displays and street-level activities (Cao et al., 2008; Reeves et al., 2005), where visible, playful activities naturally draw passersby into becoming spectators and even spontaneous participants. Similarly, during our field trials, once a small group initiated MR street play, curious bystanders consistently gathered around them. Players immersed in digital battles or performing dramatic gestures transformed ordinary streets into lively theaters of interaction. This mirrors previous findings showing how street play and public interactive technologies dynamically convert urban spaces into stages for public performance and social interaction (Stevens, 2007; Benford and Giannachi, 2011).
@@
Beyond merely drawing attention, IMRSP created significant social contagion effects. Our observations demonstrated that players felt notably less awkward engaging in playful activities when accompanied by peers. The presence of a supportive social group reduced inhibitions, echoing findings from previous interactive installations where collective engagement validated individual participation in otherwise unconventional behaviors (Reeves et al., 2005; Wouters et al., 2016). Additionally, bystanders frequently mirrored the players’ enthusiasm through gestures, applause, or informal engagement, reinforcing the playful atmosphere. This contagion effect underscores IMRSP’s potential to leverage social facilitation, effectively lowering barriers for hesitant individuals to engage in public MR interactions.
@@
Moreover, participants frequently embraced the public attention as an opportunity for deliberate showmanship (Reeves et al., 2005), exaggerating movements and reactions to entertain and involve bystanders. Rather than feeling inhibited by an audience, these players became energized, further emphasizing the performative and expressive dimensions of IMRSP. This aligns with prior research on embodied interaction and performance-led HCI approaches, where technology encourages expressive and performative engagement (Dourish, 2001; Benford et al., 2013). Such intentional performances amplified community interactions and fostered a positive feedback loop, attracting additional spectators and participants.
@@
Collectively, these findings reveal how IMRSP actively transforms public spaces into shared arenas of playful social interaction and performance. The street, traditionally governed by subtle social norms, becomes a vibrant social stage where traditional boundaries between player and audience blur. As MR street play grows more prevalent, understanding and managing these emergent social dynamics will be essential for facilitating inclusive and positive public interactions in MR-rich urban futures.
@@ 6.2. The Frictional Path to Widespread MR Street Play Adoption
6.2.1. Contrast of Social Acceptance: On-site vs. Online Perspectives
Our study revealed a notable gap in social acceptance between those who experienced IMRSP firsthand and those who observed it remotely through recorded gameplay videos. Participants who directly engaged in MR street play consistently reported highly positive experiences, describing IMRSP as enjoyable, immersive, and socially rewarding. They appreciated the embodied interaction facilitated by MR headsets, highlighting that playing alongside friends and maintaining awareness of their physical surroundings significantly reduced feelings of isolation or awkwardness. This on-site enthusiasm aligns closely with prior research emphasizing the role of collaborative experiences in alleviating social discomfort during public interactions (Reeves et al., 2005; Huggard et al., 2013; Benford et al., 2013).
In sharp contrast, remote online observers, who lacked direct interaction and immersion, expressed far more ambivalence and skepticism towards IMRSP. Online commenters frequently voiced concerns about social awkwardness or embarrassment associated with observed behaviors, often characterizing public MR play as “strange” or “inappropriate.” Safety concerns were also prominent among these remote viewers, who worried that players might become distracted, increasing the risk of accidents or collisions in public settings. These critical perceptions resonate with earlier societal reactions to novel wearable technologies such as Google Glass, where non-users perceived the technology as intrusive or socially disruptive due to unfamiliarity and lack of direct exposure (Pelikan et al., 2024; McGrane, 1994).
This perception gap underscores a critical challenge in achieving broader public acceptance of IMRSP. On-site players, immersed in the embodied and collaborative MR experience, quickly grasp its social and recreational value, seeing the technology as socially acceptable and enriching. Conversely, external observers, lacking firsthand immersion, primarily witness unconventional behaviors without the contextual understanding or immediate social engagement afforded by direct participation. As such, these observers often interpret IMRSP as violating established social norms or expectations, particularly regarding appropriate behavior in shared public spaces—a reaction aligned with Goffman’s insights on social norms and the management of impressions in public contexts (Goffman, 1966).
@@
6.2.2. Ethical, Safety, and Privacy Concerns in Public MR Play